# OctreeOcc: Efficient and Multi-Granularity Occupancy Prediction Using Octree Queries

### 09 Dec 2023 v2

@@ 1 Introduction
(a) Sparsity of 3D scene
(b) Reference of octree in 2D
(c) Difference between octree representation and grid representation
Figure 1: Illustration of the sparsity of 3D scene and octree representation. (a) shows that the semantic categories only account for a small fraction of the space, and that modeling empty regions densely affects computational efficiency. (b) and (c) demonstrate the superiority of octree representations, where we can apply different granularities for different scales of objects or spaces, which reduces computational overhead while preserving spatial information. Translucent areas represent empty voxels.
@@
Existing occupancy prediction methods [8, 19, 42, 10, 45] typically construct dense grid representations, same as the occupancy ground truth. It is a natural and straightforward solution but does not take the property of 3D scene geometry into consideration. The 3D world is composed of foreground objects and background regions with shapes and sizes varying significantly. For example, in the occupancy ground truth, there can be more than 90% empty regions as shown in Fig. 0(a). The regions occupied by objects such as buses and construction vehicles are also much larger than small objects like traffic cones. Thus, using the same resolution of voxels to represent the scene is sub-optimal. As a result, the aforementioned works struggle with large computational overhead. Other recent works attempted to mitigate the heavy memory footprint with other representations, such as 2D planes and coarse voxels exploited in TPVFormer [9] and PanoOcc [41]. VoxFormer [17] leverages the depth estimation to model non-empty regions. However, they still suffer from the loss of spatial information (i.e. 2D planes) and rely on the depth estimation quality.
@@
In contrast to previous attempts, we tackle this issue with the octree [25] representation that exactly accommodates object shapes and semantic regions with varying sizes. As a tree-based data structure, it recursively divides the 3D space into eight octants, thus allowing coarse spatial partition for large regions but fine-grained processing for small objects or complex details (Fig. 0(b) and 0(c)). Incorporating this representation into the occupancy prediction framework, we propose OctreeOcc, an efficient and multi-granularity framework using octree queries. It constructs octree queries by predicting the octree structure from the stored features of leaf nodes at each level of the tree. Such 3D structure prediction from 2D images is challenging due to the lack of depth and occlusion issues. To address this problem, we first propose Semantic-guided Octree Initialization that incorporates image semantic information to produce more accurate initial structures. To further improve the octree prediction, we devise an Iterative Structure Rectification module that predicts new octree structure from the encoded query to rectify the low-confidence region of the original prediction.
@@
Our extensive evaluations against state-of-the-art occupancy prediction methods demonstrate that OctreeOcc outperforms others on nuScenes and SemanticKITTI datasets, reducing computational overhead by $15\%-24\%$ for dense-grid representation-based methods. Ablation studies further validate the effectiveness of each module within our method. In summary, our contributions can be summarized as follows:
@@
    We introduce OctreeOcc, a 3D occupancy prediction framework based on multi-granularity octree query. The scheme achieves spatial sparsification, and effectively reduces the number of voxels required to represent the scene while preserving the essential spatial information.
@@
    We develop the semantic-guided octree initialization module and iterative structure rectification module to enable the network to have a good initialization and dynamically adjust the octree for a more effective representation.
@@ 2.1 Camera-based 3D Perception
Camera-based 3D perception has gained significant traction in recent years due to its ease of deployment, cost-effectiveness, and the preservation of intricate visual attributes. According to the view transformation paradigm, these methods can be categorized into three distinct types. LSS-based approaches[29, 8, 15, 16, 24, 28] explicitly lift multi-view image features into 3D space through depth prediction. Another category of works[18, 44, 11] implicitly derives depth information by querying from 3D to 2D. Notably, projection-free methods[22, 23, 39, 21] have recently demonstrated exceptional performance. While commendable progress has been made in detection, this approach compromises the comprehensive representation of the overall scene structure within 3D space and proves less effective in recognizing irregularly shaped objects. Consequently, there is a burgeoning interest in methodologies aimed at acquiring a dense voxel representation through the camera, facilitating a more comprehensive understanding of 3D space.
@@
3D occupancy prediction involves the prediction of both occupancy and semantic attributes for all voxels encompassed within a three-dimensional scene, particularly valuable for autonomous vehicular navigation. Recently, some valuable datasets [35, 40, 33] have been proposed, boosting more and more research works [27, 34] in this field. Most of the research in occupancy prediction focuses on dense voxel modeling. MonoScene[3] pioneers a camera-based approach using a 3D UNet architecture. OccDepth[26] improves 2D-to-3D geometric projection using stereo-depth information. OccFormer[45] decomposes the 3D processing into the local and global transformer pathways along the horizontal plane. SurroundOcc[42] achieves fine-grained results with multiscale supervision. Symphonies[10] introduces instance queries to enhance scene representation.
Nevertheless, owing to the high resolution of regular voxel representation and sparse context distribution in 3D scenes, these methods encounter substantial computational overhead and efficiency issues. Some approaches recognize this problem and attempt to address it by reducing the number of modeled voxels. For instance, TPVFormer[9] employs a strategy of modeling the three-view 2D planes and subsequently recovering 3D spatial information from them. However, its performance degrades due to the lack of 3D information. PanoOcc[41] initially represents scenes at the coarse-grained level and then upsamples them to the fine-grained level, but the lack of information from coarse-grained modeling cannot be adequately addressed by the up-sampling process. VoxFormer[17] mitigates computational complexity by initially identifying non-empty regions through depth estimation and modeling only those specific areas. However, the effectiveness of this process is heavily contingent on the accuracy of depth estimation.
In contrast, our approach provides different granularity of modeling for different regions by predicting the octree structure, which reduces the number of voxels to be modeled while preserving the spatial information, thereby reducing the computational overhead and maintaining the prediction accuracy.
@@
2.3 Octree-Based 3D Representation
The octree technique[25] involves recursively partitioning three-dimensional space into eight octants. Widely utilized in computer graphics, it is integral to applications such as rendering, shape reconstruction, and collision detection. Due to its spatial efficiency and ease of implementation on GPU architectures, researchers have applied octrees in 3D point cloud learning and various related works. Some research[37, 38, 13] leverages the octree structure to efficiently represent and analyze the shape of the point cloud. OctFormer[36] and OcTr [46] utilize octree multi-granularity features to capture a comprehensive global context, thereby enhancing the efficiency of understanding point clouds at the scene level. Furthermore, certain studies [6, 30] adopt octree representation for effectively compressing point cloud data. These instances highlight the versatility and effectiveness of octree-based methodologies in point cloud analysis and processing applications. However, unlike constructing an octree from a 3D point cloud, predicting the octree structure of a 3D scene from images is a challenging task owing to the absence of explicit spatial information inherent in 2D representations. In order to address this complexity, we take the semantic information as a precursor to initializing the octree structure. Subsequently, we iteratively update the structure, ensuring a continual rectification that aligns more accurately with the scene.
3 Methods
3.1 Problem Steup
Camera-based occupancy prediction aims to predict the present occupancy state and semantics of each voxel grid within the scene using input from multi-view camera images. Specifically, we consider a set of $N$ multi-view images $I$ = $\{I_{i}\in\mathbb{R}^{3\times H\times W}\}_{i=1}^{N}$, together with camera intrinsics $K$ = $\{K_{i}\in\mathbb{R}^{3\times 3}\}_{i=1}^{N}$ and extrinsics $T$ = $\{T_{i}\in\mathbb{R}^{4\times 4}\}_{i=1}^{N}$ as input and the objective of the model is to predict the 3D occupancy $O\in\mathbb{R}^{X\times Y\times Z\times K}$, where $X$, $Y$, $Z$ denote the volume resolution, and $K$ represents the number of semantic categories.
3.2 Overall Architecture
As illustrated in Figure 2, OctreeOcc primarily comprises four modules. Initially, the image backbone extracts multi-scale features from multi-view images. Then we randomly initialize the dense occupancy queries $Q_{dense}\in\mathbb{R}^{X\times Y\times Z\times C}$, where $X$, $Y$, $Z$ are the spatial shape of occupancy queries, and C is the channel dimension. Subsequently, leveraging an image segmentation prior (Section 3.4), the octree initialization module transforms $Q_{dense}$ into sparse octree queries $Q_{octree}$ (Section 3.3). Facilitated by the octree encoder, image features are conveyed to the octree query, and the octree structure undergoes continuous refinement (Section 3.5). Finally, an octree decoder translates octree queries back into dense grid queries, culminating in the generation of outputs for the occupancy prediction task.
@@ 3.3 Octree Query
Given that objects within a 3D scene exhibit varying granularities, employing dense queries for processing coarse-grained objects leads to performance wastage and inefficiency. To address this issue, we introduce sparse and multi-granularity octree queries, leveraging the octree structure. This approach enables the creation of flexible voxel representations tailored to semantic regions across different scales. To construct an octree query, we initiate the process by predicting the octree mask. This mask serves as a critical element in acquiring the octree structure, facilitating the conversion of $Q_{dense}$ into $Q_{octree}$. Denoted as $M_{o}\in\mathbb{R}^{\frac{X}{2^{l}},\frac{Y}{2^{l}},\frac{Z}{2^{l}}}$, where $l=1,2,\ldots,L-1$, the octree mask signifies the likelihood that each voxel at every level undergoes a split into octants. Here, $L$ represents the depth of the octree.
Based on $M_{o}$ and a query selection ratio denoted as $\alpha=\alpha^{1},\alpha^{2},\ldots,\alpha^{L-1}$, which governs the number of subdivisions at each level, we can define the hierarchical structure of the octree, specifically determining how queries are assigned to respective levels. The procedure is as follows. For level 1, queries with $M_{o}^{1}$ values within the top $\alpha^{1}$ percentile are identified as candidates for octant subdivision. Subsequently, within octants that have undergone one previous subdivision at the next level, queries are once again selected based on their values falling within the top $\alpha^{2}$ percentile, initiating another round of subdivision. This process continues iteratively until reaching the final level of the octree.
By leveraging structural information, we can transform dense query into sparse and multi-granularity octree query. Specifically, we downsample the $Q_{dense}$ at each level using average pooling. Within each level, we selectively retain queries that have no need for further subdivision into octants (i.e., leaf queries). By storing all the constructed leaf queries, we form the $Q_{octree}\in\mathbb{R}^{N\times C}$, where $N=N_{1}+N_{2}+\ldots+N_{L}$ denotes the number of leaf queries. This information is updated in the octree encoder module along with the octree mask.
At the same time, leveraging the octree structure allows for the straightforward conversion of sparse octree queries into dense queries to match the desired output shape. Specifically, for a query $q_{octree}$ at level $l$ with the index $(a,b,c)$, its corresponding $8^{(}L-l)$ children nodes’ indexes in level $L$ are given by $(a\times 2^{L-l}+a_{offset},b\times 2^{L-l}+b_{offset},c\times 2^{L-l}+c_{offset})$, where $a_{offset},b_{offset},c_{offset}=0,1,\ldots,2^{L-l}$ and L denotes the depth of the octree. In this process, we assign $q_{octree}$’s feature to all of these positions. By repeating this procedure for all queries at each level, we effectively convert the $Q_{octree}$ into $Q_{dense}$.
@@ 3.4 Semantic-Guided Octree Initialization
Predicting octree structure from an initialised query via neural network can yield sub-optimal results due to the inherent lack of meaningful information in the query. To address this limitation, we employ the image prior as a pivotal reference point. Specifically, we adopt UNet[32] to segment the input multi-view images $I$ and obtain the segment map $I_{seg}\in\mathbb{R}^{H\times W\times K_{seg}}$. We then generate sampling points $\{p_{i}=(x_{i},y_{i},z_{i}),i\in X\times Y\times Z\}$ corresponding to a single occupancy query $Q^{p_{i}}$ at grid region center $p_{i}=(x_{i},y_{i},z_{i})$, and then project these points to different image views. The projection between sampling points $p_{i}$ and its corresponding 2D reference point $(u_{ij},v_{ij})$ on the $j$-th image view is formulated as:
$\pi_{j}(p_{i})=(u_{ij},v_{ij})$ (1)
where $\pi_{j}(p_{i})$ denotes the projection of the $i$-th sampling point at location $p_{i}$ on the $j$-th camera view.
We project the point $p_{i}$ onto the acquired semantic segmentation map $I_{seg}$ through the described projection process. To address the impact of the long-tailed distribution inherent in the dataset, we assign varying weights to voxels projected onto different semantic categories. After obtaining the initial confidence for each voxel, we downsample it to different octree levels to formulate the initial octree mask. This mask will be further refined in the octree encoder.

### 19 Mar 2024 v3

@@ Abstract
Keywords:
3D Scene Understanding Occupancy Prediction Octree
@@ 1 Introduction
(a) Instance Size Inbalance
(b) 3D and 2D Visualization of Octree Structures
Figure 1: Scale difference of various categories and octree representation. (a) presents a comparison of the average space occupied by each type of object, indicating different granularities are required for different semantic regions. (b) shows the superiority of octree representations, where we can apply specific granularity not only for different objects but also for different parts of the object, which can reduce computational overhead while preserving spatial information.
@@
Existing occupancy prediction methods [9, 22, 51, 12, 56] typically construct dense and regular grid representations, same as the ground truth. While such approach is intuitive and direct, it overlooks the statistical and geometric properties of 3D environments. In fact, the 3D scene is composed of foreground objects and background regions with various shapes and sizes. For example, the space occupied by larger objects, such as buses, are considerably more extensive than that taken up by smaller items like traffic cones (Fig. 0(a)). Consequently, employing a uniform voxel resolution to depict the scene proves to be inefficient, leading to computational waste for larger objects and a lack of geometry details for smaller ones. Considering the large computation cost of aforementioned works, some recent works attempted to mitigate the heavy memory footprint by utilizing other representations, such as 2D planes in TPVFormer [11] and coarse voxels in PanoOcc [50], or modeling non-empty regions by depth estimation [18]. However, these methods suffer from the loss of spatial information because of too coarse representations or accumulated estimated depth errors.
@@
To reduce the computational overhead and meanwhile improve the prediction accuracy, we propose to use octree [29] representation for the occupancy prediction, which can flexibly adapt to objects and semantic regions with various shapes and sizes. As a tree-based data structure, it recursively divides the 3D space into eight octants, thus allowing coarse spatial partition for large regions and fine-grained processing for small objects or complex details (Fig. 0(b)). Incorporating octree representation, we propose OctreeOcc, an efficient and multi-granularity method for occupancy prediction. It constructs octree queries by predicting the octree structure from the stored features of leaf nodes at each level of the tree. However, directly predicting 3D structure from 2D images is challenging due to the lack of depth and occlusion issues. To address this problem, we first propose Semantic-guided Octree Initialization that incorporates image semantic information to produce more accurate initial structures. And then, we devise an Iterative Structure Rectification module that predicts new octree structure from the encoded query to rectify the low-confidence region of the original prediction, further improving the prediction precision.
@@
Our extensive evaluations against state-of-the-art occupancy prediction methods show that OctreeOcc outperforms others on nuScenes and SemanticKITTI datasets, reducing computational overhead by $15\%-24\%$ for dense-grid-based methods. Ablation studies further validate the effectiveness of each module within our method. Our contributions can be summarized as follows:
@@
    We introduce OctreeOcc, a 3D occupancy prediction approach based on multi-granularity octree queries. This method facilitates spatial sparsification, significantly decreasing the number of voxels needed to accurately depict a scene, yet retains critical spatial details.
@@
    We propose a semantic-guided octree initialization module and and iterative structure rectification module. These enhancements empower the network with a robust initial setup and the ability to dynamically refine the octree, leading to a more efficient and effective representation.
@@ 2.1 Camera-based 3D Perception
Camera-based 3D perception has gained significant traction in recent years due to its ease of deployment, cost-effectiveness, and the preservation of intricate visual attributes. According to the view transformation paradigm, these methods can be categorized into three distinct types. LSS-based approaches[34, 9, 19, 20, 27, 33] explicitly lift multi-view image features into 3D space through depth prediction. Another category of works[21, 53, 13] implicitly derives depth information by querying from 3D to 2D. Notably, projection-free methods[25, 26, 48, 24] have recently demonstrated exceptional performance. While commendable progress has been made in detection, this approach compromises the comprehensive representation of the overall scene in 3D space and proves less effective in recognizing irregularly shaped objects. Consequently, there is a burgeoning interest in methodologies aimed at acquiring a dense voxel representation through the camera, facilitating a more comprehensive understanding of 3D space.
@@
3D occupancy prediction involves the prediction of both occupancy and semantic attributes for all voxels encompassed within a three-dimensional scene, particularly valuable for autonomous vehicular navigation. Recently, some valuable datasets [43, 49, 39] have been proposed, boosting more and more research works [32, 40, 3, 30, 56, 51, 12, 11, 50, 18, 10, 28, 54, 31, 55, 38] in this field. Most of the research in occupancy prediction focuses on dense voxel modeling. MonoScene[3] pioneers a camera-based approach using a 3D UNet architecture. OccDepth[30] improves 2D-to-3D geometric projection using stereo-depth information. OccFormer[56] decomposes the 3D processing into the local and global transformer pathways along the horizontal plane. SurroundOcc[51] achieves fine-grained results with multiscale supervision. Symphonies[12] introduces instance queries to enhance scene representation.
Nevertheless, owing to the high resolution of regular voxel representation and sparse context distribution in 3D scenes, these methods encounter substantial computational overhead and efficiency issues. Some approaches recognize this problem and attempt to address it by reducing the number of modeled voxels. For instance, TPVFormer[11] employs a strategy of modeling the three-view 2D planes and subsequently recovering 3D spatial information from them. However, its performance degrades due to the lack of 3D information. PanoOcc[50] initially represents scenes at the coarse-grained level and then upsamples them to the fine-grained level, but the lack of information from coarse-grained modeling cannot be adequately addressed by the up-sampling process. VoxFormer[18] mitigates computational complexity by initially identifying non-empty regions through depth estimation and modeling only those specific areas. However, the effectiveness of this process is heavily contingent on the accuracy of depth estimation. In contrast, our approach provides different granularity of modeling for different regions by predicting the octree structure, which reduces the number of voxels to be modeled while preserving the spatial information, thereby reducing the computational overhead and maintaining the prediction accuracy.
2.3 Octree-Based 3D Representation
The octree structure[29] has been widely used in the field of computer graphics to boost the process of rendering or reconstruction[7, 15, 41, 46]. Due to its spatial efficiency and ease of implementation on GPU architectures, researchers have also applied octrees in efficient point cloud learning and various related works [45, 47, 16, 42]. OctFormer[44] and OcTr [57] utilize multi-granularity features of octree to capture a comprehensive global context, thereby enhancing the efficiency of understanding point clouds at the scene level. Furthermore, certain studies [6, 35] adopt octree representation for effectively compressing point cloud data. These works have highlighted the versatility and effectiveness of octree-based methodologies in point cloud analysis and processing applications. However, unlike constructing an octree from 3D point clouds, we are the first to predict the octree structure of a 3D scene from images, which is more challenging owing to the absence of explicit spatial information inherent in 2D images.
3 Methodology
@@
In this section, we introduce details of our efficient and multi-granularity occupancy prediction method OctreeOcc. After defining the problem and providing an overview of our method in Sec. 3.1 and 3.2, we introduce key components of our method in order. In Sec. 3.3, we outline how we define octree queries and transform dense queries into octree queries. Next, we utilize image semantic priors to construct a high-quality initialized octree structure, as detailed in Sec. 3.4. Once the initialized octree query is obtained, we encode it and refine the octree structure in Sec. 3.5. Finally, Sec. 3.6 and 3.7 describe the process of generating octree ground truth and how to supervise the network, respectively.
3.1 Problem Setup
Camera-based occupancy prediction aims to predict the present occupancy state and semantics of each voxel grid within the scene using input from multi-view camera images. Specifically, we consider a set of $N$ multi-view images $I$ = $\{I_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$, together with camera intrinsics $K$ = $\{K_{i}\in\mathbb{R}^{3\times 3}\}_{i=1}^{N}$ and extrinsics $T$ = $\{T_{i}\in\mathbb{R}^{4\times 4}\}_{i=1}^{N}$ as input, and the objective of the model is to predict the 3D semantic voxel volume $O\in\{w_{0},w_{1},...,w_{C}\}^{X\times Y\times Z}$, where $H$, $W$ indicate the resolution of input image and $X$, $Y$, $Z$ denote the volume resolution (e.g. 200$\times$200$\times$16). The primary focus lies in accurately distinguishing the empty class ($w_{0}$) and other semantic classes ($w_{1}\sim{w_{C}}$) for every position in the 3D space, which entails the network learning both the geometric and semantic information inherent in the data.
3.2 Overview
Given a set of multi-view images $I=\{I_{i}\in\mathbb{R}^{H\times W\times 3}\}_{i=1}^{N}$, we extract multi-view image features $\mathbb{F}=\{F_{i}\in\mathbb{R}^{H\times W\times C}\}_{i=1}^{N}$. Simultaneously, we initialize the dense voxel query $Q_{dense}\in\mathbb{R}^{X\times Y\times Z\times C}$. To enhance computational efficiency, we transform $Q_{dense}$ into sparse representation $Q_{octree}\in\mathbb{R}^{N\times C}$, leveraging octree structure information ($i.e.$ octree mask) derived from segmentation priors. During encoding, we utilize $Q_{octree}$ to gather information, including temporal fusion and sampling from image features $\mathbb{F}$, while also rectifying the octree structure. Upon encoding $Q_{octree}$, to conform to the output format, we convert it back to $Q_{dense}$ and apply a Multi-Layer Perceptron (MLP) to obtain the final occupancy prediction $O\in\mathbb{R}^{X\times Y\times Z\times K}$. Here, $H$, $W$ indicate the resolution of input image and $X$, $Y$, $Z$ denote the volume resolution. $N$ means the number of octree query ,$C$ denotes the feature dimension and $K$ indicates the number of classes.
@@ 3.3 Octree Query
Given that objects within 3D scenes exhibit diverse granularities, employing dense queries [56, 51] overlooks this variation and leads to inefficiency. To address this, we propose sparse and multi-granularity octree queries, leveraging the octree structure. This approach creates adaptable voxel representations tailored to semantic regions at different scales.
Octree Mask. To effectively construct the octree query, it’s essential to understand its underlying structure. An octree partitions each node into eight child nodes within 3D space, each representing equal subdivisions of the parent node. This recursive process begins with the initial level and proceeds with gradual splitting. At every level, a voxel query serves either as a leaf query if it remains unsplit or as a parent query if it undergoes division. We obtain this geometric information by maintaining a learnable octree mask, denoting as $M_{o}=\{M_{o}^{l}\in\mathbb{R}^{\frac{X}{2^{l}},\frac{Y}{2^{l}},\frac{Z}{2^{l}}}\}_{l=1}^{L-1}$, where $X$, $Y$, $Z$ denote the ground truth’s volume resolution. The $L$ denotes the depth of the octree, representing the existence of $L$ different resolutions of queries, with $L-1$ splits being performed from the top to the bottom of the octree. The value in the octree mask represents the probability that a voxel at that level requires a split, which is initialized through segmentation priors (Sec. 3.4), rectified during query encoding (Sec. 3.5), and supervised by octree ground truth (Sec. 3.7).
Transformation between octree query and dense voxel query. During the query encoding process, we prioritize efficiency by leveraging octree queries. This involves transforming the initial dense queries $Q_{dense}$ into octree queries $Q_{octree}$ using learned structural information. To determine the octree structure, we need to binarise the learned octree mask. Due to the fact that most of the voxels in the scene at various resolutions do not necessitate splitting, neural network-based prediction binarization is susceptible to pattern collapse, tending to predict all as non-split, leading to a decrease in performance. To mitigate this issue, we introduce a manually defined query selection ratio, where a subset of voxels with the highest probability of splitting is selected for division through the top-k mechanism.
The transformation from $Q_{dense}$ to $Q_{octree}$ begins at the finest granularity, we downsample $Q_{dense}$ to each level through average pooling and retain queries that do not require splitting (leaf queries) with the assistance of the binarized octree mask. This process iterates until reaching the top of the octree. By retaining all leaf queries, we establish $Q_{octree}\in\mathbb{R}^{N\times C}$, where $N=N_{1}+N_{2}+\ldots+N_{L}$ represents the total count of leaf queries, L indicates the depth of octree. Conversely, applying the inverse operation of this process allows the conversion of $Q_{octree}$ back into $Q_{dense}$ for the final output. Further details are provided in Supplementary.
@@ 3.4 Semantic-Guided Octree Initialization
Predicting octree structure from an initialised query via neural network can yield sub-optimal results due to the inherent lack of meaningful information in the query. To overcome this limitation, we employ the semantic priors inherent in images as crucial references. Specifically, we adopt UNet[37] to segment the input multi-view images $I$ and obtain the segment map $I_{seg}=\{I_{seg}^{i}\in\mathbb{R}^{H\times W}\}_{i=1}^{N}$. We then generate sampling points $p=\{p_{i}\in\mathbb{R}^{3}\}_{i=1}^{X\times Y\times Z}$, with each point corresponding to the center coordinates of dense voxel queries. Subsequently, we project these points onto various image views. The projection from sampling point $p_{i}=(x_{i},y_{i},z_{i})$ to its corresponding 2D reference point $(u_{ij},v_{ij})$ on the $j$-th image view is formulated as:
$\pi_{j}(p_{i})=(u_{ij},v_{ij}),$ (1)
where $\pi_{j}(p_{i})$ denotes the projection of the $i$-th sampling point at location $p_{i}$ on the $j$-th camera view. We project the point $p_{i}$ onto the acquired semantic segmentation map $I_{seg}$ through the described projection process. To prioritize effective areas like foreground objects and buildings in the initial structure, we employ an unbalanced assignment approach, where the highest weight is assigned to voxels projecting to foreground areas, followed by lesser weights for voxels projecting to buildings or vegetation, and the least weight for voxels projecting to roads, etc.
During this process, we determine the weights of each voxel at the finest granularity, denoted as $W\in\mathbb{R}^{X\times Y\times Z}$. Subsequently, we employ average pooling to downsample $W$ to each level of the octree, resulting in an initial octree mask $M_{o}=\{{M_{o}^{l}\in\mathbb{R}^{\frac{X}{2^{l}},\frac{Y}{2^{l}},\frac{Z}{2^{l}}}}\}_{l=1}^{L-1}$. Here, $X$, $Y$, and $Z$ represent the resolution of the ground truth volume, while $L$ denotes the depth of the octree. Further details are provided in the Supplementary.